OpenStudy (anonymous):

list three different ways to write 5^11 as the product of two powers. explain why all there of your expressions are equal to 5^11

hero (hero):

\(5^3*5^8 = 5^{3 + 8} = 5^{11}\) is one way to do it. Can you think of another way @princess_51 ?

OpenStudy (anonymous):

nope

hero (hero):

\(5^2*5^9 = 5^{2 + 9} = 5^{11}\) is one other way to do it. Keep in mind, the rule of exponents related to this is \(a^{b}*a^{c} = a^{b + c}\) In this case, \(a = 5\) and the values for \(b\) and \(c\) must add to get eleven.

OpenStudy (anonymous):

one more please

OpenStudy (bee_see):

Just ask yourself what two numbers add up to 11.